# Richelieu: Self-Evolving LLM-Based Agents for AI Diplomacy

### v2

@@ 5 Experiment
In the experiments, our goal is to answer the following questions: 1) Mastery of Non-Press Diplomacy: Can our agent master the non-press diplomacy against baselines? 2) Competing with State-of-the-Art: Can our agent surpass the performance of the current state-of-the-art agents in press diplomacy? 3) Compatibility with LLMs: Can our self-evolving framework be compatible with different LLMs? 4) Contribution of Framework Modules: Do the individual modules within our framework contribute to the overall improvement of our agent’s performance? 5) Social Reasoning: Can Richelieu accurately infer the true intentions of other players and reasonably determine the relationships of ally or enemy with them?
@@ 5.1 Experimental Setup
Environment. The widely-used open source Diplomacy game platform introduced by (Paquette et al., 2019) is adopted for evaluating Richelieu against other models. It is easy to switch between no-press (with communication/negotiation between players) and press (no communication between players) games based on this platform, facilitating comparison on both settings. The platform also contains over 10,000 human game data on which previous approaches are trained. Note that our method does not need them. In each game, a model will play the role of one randomly selected country to compete against countries controlled by other methods. It wins if occupying all the supply centers and loses vice versa.
@@
Evaluation Metrics. We evaluate the models based on the results of multiple rounds of games. In each round, the model is randomly assigned a country to control. Typically, 1000 rounds are played to obtain the average results. We evaluate the models in two metrics. One is based on the win rate, Most SC rate, survived rate, and defeated rate. There are four possible outcomes for each country in the game. If a country loses all its supply centers (SC), it is eliminated and recorded as “defeated". If a country occupies 18 or more out of 34 supply centers, the game ends, and that country is recorded as “win", while other countries are recorded as “defeated". In other cases, the game ends in a draw. The country with the most supply centers is recorded as “Most SC", the countries that have been eliminated are recorded as “defeated", and the other countries are recorded as “Survived". The other is based on the scores obtained by the models after multiple rounds of competition. To compare the capabilities of multiple models, we use C-Diplo Argir(Archer, 2024), a scoring system. This system is used in many international diplomacy competitions. The scoring method is as follows: If a player wins by occupying 18 or more supply centers, the player scores 93 points, and each of the other six players scores 1 point. If the game ends in a draw, the player with the most centers scores 37 points. The second player with the most centers scores 14 points. The third player with the most centers scores 7 points. Each player scores 1 point per center owned. Each player also scores 1 point for participating. In this way, regardless of the game outcome, a total of 99 points will be distributed among the players in each game.
Baselines. We select six previous models as baselines for comparison. Among them, Cicero(Bakhtin et al., 2022) by Meta is a diplomacy model with a negotiation module. The other five are no-press diplomacy models, including the SL-DipNet and RL-DipNet (Paquette et al., 2019), the BRPI (Anthony et al., 2020), the SearchBot (Gray et al., 2020), and the DORA(Bakhtin et al., 2021). We also build a LLM-based agent, AutoGPT Yang et al. (2023a). In experiments, we set a temperature of 0.3 to ensure a relatively stable generation of LLM policies. The overall reasoning framework also ensure the stability and consistency in the AI agent’s performance.
@@ 5.2 Results
Massively Play with Baselines on no-press setting. We let Richelieu compete with the other six models including Cicero(Bakhtin et al., 2022), SL-DipNet and RL-DipNet (Paquette et al., 2019), BRPI (Anthony et al., 2020), SearchBot (Gray et al., 2020), and DORA(Bakhtin et al., 2021) on No-Press Diplomacy, in which players make moves without communication. Figure  4 indicates that Richelieu outperforms other previous models relying on human game data. In contrast, Richelieu does not need such data but outperforms these methods by a clear margin, which demonstrates the outstanding planning capability of Richelieu.
@@
Generalization of self-evolving framework to diverse LLMs.
Llamma 3 To demonstrate the effectiveness of our framework in a variety of LLM, we conducted experiments using four models: GPT4.0, ERNIE Bot, Spark Desk, and Llamma 3. As the number of training iterations increases, Richelieu’s win rate steadily improves while the defeated rate declines, ultimately reaching a relatively stable outcome. This suggests that our self-play method is effective. After training, the win rate using GPT4.0 increased from 1.5% lower than Cicero’s to about 0.7% higher than Cicero’s. The win rate using llama3 increased from 2.3% lower than Cicero’s to almost equal to Cicero’s. The win rates using Models Spark Desk and ERNIE Bot increased from 3% and 4% lower than Cicero’s to 0.7% and 1.6% lower than Cicero’s, respectively. The experimental results show that, despite variations in Richelieu’s performance due to the inherent differences in the capabilities of these LLMs, as illustrated in Figure 5, our framework and training approach significantly enhance the capabilities of all large language models.This indicates the generalization of a self-evolving framework to various LLMs.
@@ 7 Limitations and Future Work
Our framework is capable of applying to most social interaction tasks. Most components in our framework can be easily generalized to a new task by modifying the content. Social reasoning enables the agent to handle complex and dynamic social relationships. The negotiation pipeline opens the potential of communicating with others to prob the other’s mind or reach a consensus. The hierarchical strategy with reflection enhances the ability to handle long-term planning. The self-evolving mechanism (reflection with self-play memory) further improves the overall performance without manual supervision. These modules cover most of the challenges in multi-agent interactions. The potential applications of such an AI agent are vast, ranging from simulated diplomatic environments to real-world assistance and analysis. In future research, we intend to develop a more realistic game space, characterized by incomplete information and multi-player games, to enhance and refine our model further. We will also extend the framework to other multi-agent scenarios, including embodied interactions (Zhong et al., 2023; Ci et al., 2023; Chen et al., 2023), sensor networks (Wang et al., 2021; Xu et al., 2020; Pan et al., 2022; Li et al., 2020), and video games (Wang et al., 2024a; Ma et al., 2024). This framework can also be employed to develop various applications. For instance, in the fields of economics and finance, we intend to utilize it to create business analytics and negotiation models.
Code Release
Implementation of the proposed method can be found at this link for reproducing our results: https://github.com/todexter3/Richelieu.git
@@ Acknowledgements
This work was supported by the National Science and Technology Major Project (MOST-2022ZD0114900), NSF 6240073415, and Qualcomm University Research Grant.
@@ B.1 Cases of the Effect of the Memory from Self-Playing and Collaboration
After self-play, using the historical experience from the memory module, Richelieu adjusted his strategy. Richelieu foresees Russia becoming the most threatening enemy in the future and sets a sub-goal of weakening Russia, ally with Austria and Turkey, and attack Britain. Figure 6(b) shows the state and the negotiations after self-play, where we actively sought an armistice alliance with Austria to make Austria concentrate their forces against the Russian attack. In the subsequent negotiation phase, Richelieu proactively proposes ending the war with Austria, despite holding an advantage in this conflict. Richelieu promises Austria that if it ceases hostilities and attacks Russia, Richelieu will assist Austria in defending against any attacks from England. The negotiations are successful. Austria accepted Richelieu’s proposal, and the two countries reached an agreement to exchange the supply centers of Napoli and Munich. During the action phase, Austria moves its troops from Venice to Apulia in preparation for capturing Napoli in the next turn, while the rest of its forces are repositioned to the eastern regions bordering Russia to defend against Russian attacks and compete for supply centers. French units occupy Munich and prepare to advance on Russian territories such as Berlin. Meanwhile, French units support Austria in the Holland and Belgium regions. In this round, we ultimately achieved a better result——Most SC. This is also a great example that highlights our model’s ability to collaborate effectively with other players.
@@ B.2 Case of Avoiding Deception
As shown in Figure  7, Richelieu controls Germany. During the negotiation phase, England proposed a ceasefire to Germany and invited Germany to form an alliance to jointly attack France. England hoped to cease the war with Germany in Holland and Belgium. Subsequently, German units would support England in attacking Brest, and then England would utilize its fleets to assist Germany in attacking Spain and Portugal. Richelieu suspected that England was deceiving Germany, as England was likely to attack territories in the north such as Belgium and Berlin after German units were redirected to support Brest. Therefore, we pretended to accept England’s alliance proposal during the negotiation process. However, at the same time, we sought out France and expressed our willingness to cease hostilities, allowing France to focus entirely on defending against England’s attacks. In the action phase, England’s actions confirmed Richelieu’s suspicions. England attacked Belgium from Holland, but because Richelieu didn’t move units in Belgium, England’s attack failed.
@@ Appendix D Ethical Consideration
The method proposed in this work has the potential for positive uses like enabling AI agents to emerge cooperation via negotiation or avoiding being fooled by fake promises (or helping human to defend against such cases). However, negative cases can also arise if the technique is used for possible fraud activities. Fortunately, there is research [Bakhtin et al., 2019][Zellers et al., 2019] dealing with such scenarios. And we also urge for more research efforts in this field to foster safe applications of similar technologies.
@@ 8. Answer: [Yes]
  9. Justification: The paper discussed the limitations of the work performed by the authors in the appendix.
@@ 13. Answer: [Yes]
  14. Justification: The paper provides the full set of assumptions and a complete (and correct) proof for each theoretical result in the method and experiment sections.
@@ 18. Answer: [Yes]
  19. Justification: The paper fully discloses all the information needed to reproduce the main experimental results in the main text experiment section and appendix section “implementation details”, “code release”.
@@ 23. Answer: [Yes]
  24. Justification: The paper provides open access to the data and code at anonymous repo provided in the appendix section “code release”.
@@ 28. Answer: [Yes]
  29. Justification: The paper specify all the training and test details to train the model in the experiment section and implementation details section of appendix.
@@ 33. Answer: [Yes]
  34. Justification: The paper reports error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiment in the experiment section.
@@ 48. Answer: [Yes]
  49. Justification: The paper discuss both potential positive societal impacts and negative societal impacts of the work performed.
@@ 53. Answer: [Yes]
  54. Justification: The paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse.
@@ 63. Answer: [Yes]
  64. Justification: New assets introduced in the paper are well documented and the documentation is provided.
@@ 68. Answer: [Yes]
  69. Justification: the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation.
@@ 73. Answer: [Yes]
  74. Justification: The paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals.

### v3

@@ 5 Experiment
In the experiments, our goal is to answer the following questions: 1) Mastery of Non-Press Diplomacy: Can our agent master the non-press diplomacy against baselines? 2) Competing with State-of-the-Art: Can our agent surpass the performance of the current state-of-the-art agents in press diplomacy? 3) Compatibility with LLMs: Can our self-evolving framework be compatible with different LLMs? 4) Contribution of Framework Modules: Do the individual modules within our framework contribute to the overall improvement of our agent’s performance? 5) Social Reasoning: Can Richelieu accurately infer the true intentions of other players and reasonably determine the relationships of ally or enemy with them? The implementation of our method can be found at: https://github.com/todexter3/Richelieu.git
@@ 5.1 Experimental Setup
Environment. The widely-used open source Diplomacy game platform introduced by Paquette et al. (2019) is adopted for evaluating Richelieu against other models. It is easy to switch between no-press (with negotiation between players) and press (no negotiation between players) games based on this platform, facilitating comparison on both settings. The platform also contains over 10,000 human game data on which previous approaches are trained. Note that our method does not need them. In each game, a model will play the role of one randomly selected country to compete against countries controlled by other methods. It wins if occupying all the supply centers and loses vice versa.
@@
[…]
Evaluation Metrics. We evaluate the models based on the results of multiple rounds of games. In each round, the model is randomly assigned a country to control. Typically, 1000 rounds are played to obtain the average results. We evaluate the models in two metrics. One is based on the win rate, Most SC rate, survived rate, and defeated rate. There are four possible outcomes for each country in the game. If a country loses all its supply centers (SC), it is eliminated and recorded as “defeated". If a country occupies 18 or more out of 34 supply centers, the game ends, and that country is recorded as “win", while other countries are recorded as “defeated". In other cases, the game ends in a draw. The country with the most supply centers is recorded as “Most SC", the countries that have been eliminated are recorded as “defeated", and the other countries are recorded as “Survived". The other is based on the scores obtained by the models after multiple rounds of competition. To compare the capabilities of multiple models, we use C-Diplo ArgirArcher (2024), a scoring system. This system is used in many international diplomacy competitions. The scoring method is as follows: If a player wins by occupying 18 or more supply centers, the player scores 93 points, and each of the other six players scores 1 point. If the game ends in a draw, the player with the most centers scores 37 points. The second player with the most centers scores 14 points. The third player with the most centers scores 7 points. Each player scores 1 point per center owned. Each player also scores 1 point for participating. In this way, regardless of the game outcome, a total of 99 points will be distributed among the players in each game.
Baselines. We select six previous models as baselines for comparison. Among them, CiceroBakhtin et al. (2022) by Meta is a diplomacy model with a negotiation module. The SL-DipNet and RL-DipNet Paquette et al. (2019), the BRPI Anthony et al. (2020), the SearchBot Gray et al. (2020), and the DORABakhtin et al. (2021) are no-press diplomacy models. We also build a LLM-based agent, AutoGPT Yang et al. (2023a). In experiments, we set a temperature of 0.3 to ensure a relatively stable generation of LLM policies. The overall reasoning framework also ensure the stability and consistency in the AI agent’s performance.
@@ 5.2 Results
Massively Play with Baselines on no-press setting. We let Richelieu compete with the other six models including CiceroBakhtin et al. (2022), SL-DipNet and RL-DipNet Paquette et al. (2019), BRPI Anthony et al. (2020), SearchBot Gray et al. (2020), and DORABakhtin et al. (2021) on No-Press Diplomacy, in which players make moves without communication. Figure  4 indicates that Richelieu outperforms other previous models relying on human game data. In contrast, Richelieu does not need such data but outperforms these methods by a clear margin, which demonstrates the outstanding planning capability of Richelieu.
@@
Generalization of self-evolving framework to different LLMs. To demonstrate the effectiveness of our framework in a variety of LLM, we conducted experiments using four models: GPT4.0, ERNIE Bot, Spark Desk, and Llama 3. As the number of training iterations increases, Richelieu’s win rate steadily improves while the defeated rate declines, ultimately reaching a relatively stable outcome. This suggests that our self-play method is effective. After training, the win rate using GPT4.0 increased from 1.5% lower than Cicero’s to about 0.7% higher than Cicero’s. The win rate using llama3 increased from 2.3% lower than Cicero’s to almost equal to Cicero’s. The win rates using Models Spark Desk and ERNIE Bot increased from 3% and 4% lower than Cicero’s to 0.7% and 1.6% lower than Cicero’s, respectively. The experimental results show that, despite variations in Richelieu’s performance due to the inherent differences in the capabilities of these LLMs, as illustrated in Figure 5, our framework and training approach significantly enhance the capabilities of all LLMs.This indicates the generalization of a self-evolving framework to various LLMs.
@@ 7 Limitations and Future Work
Our framework is capable of applying to most social interaction tasks. Most components in our framework can be easily generalized to a new task by modifying the content. Social reasoning enables the agent to handle complex and dynamic social relationships. The negotiation pipeline opens the potential of communicating with others to prob the other’s mind or reach a consensus. The hierarchical strategy with reflection enhances the ability to handle long-term planning. The self-evolving mechanism (reflection with self-play memory) further improves the overall performance without manual supervision. These modules cover most of the challenges in multi-agent interactions. The potential applications of such an AI agent are vast, ranging from simulated diplomatic environments to real-world assistance and analysis. In future research, we intend to develop a more realistic game space, characterized by incomplete information and multi-player games, to enhance and refine our model further. We will also extend the framework to other multi-agent scenarios, including embodied interactions Zhong et al. (2023); Ci et al. (2023); Chen et al. (2023), sensor networks Wang et al. (2022b); Xu et al. (2020); Pan et al. (2022); Li et al. (2020), and video games Wang et al. (2024a); Ma et al. (2024). This framework can also be employed to develop various applications. For instance, in the fields of business and finance, we intend to utilize it to create analytics and negotiation models.
@@ Acknowledgements
This work was supported by the National Science and Technology Major Project (2022ZD0114904), NSFC-6247070125, NSFC-62406034, NSFC-62406010, the State Key Lab of General Artificial Intelligence at Peking University, and Qualcomm University Research Grant.
@@ B.1 Cases of the Effect of the Memory from Self-Playing and Collaboration
After self-play, using the historical experience from the memory module, Richelieu adjusted his strategy. Richelieu foresees Russia becoming the most threatening enemy in the future and sets a sub-goal of weakening Russia, allying with Austria and Turkey, and attacking Britain. Figure 6(b) shows the state and the negotiations after self-play, where we actively sought an armistice alliance with Austria to make Austria concentrate their forces against the Russian attack. In the subsequent negotiation phase, Richelieu proactively proposes ending the war with Austria, despite holding an advantage in this conflict. Richelieu promises Austria that if it ceases hostilities and attacks Russia, Richelieu will assist Austria in defending against any attacks from England. The negotiations are successful. Austria accepted Richelieu’s proposal, and the two countries reached an agreement to exchange the supply centers of Napoli and Munich. During the action phase, Austria moves its troops from Venice to Apulia in preparation for capturing Napoli in the next turn, while the rest of its forces are repositioned to the eastern regions bordering Russia to defend against Russian attacks and compete for supply centers. French units occupy Munich and prepare to advance on Russian territories such as Berlin. Meanwhile, French units support Austria in the Holland and Belgium regions. In this round, we ultimately achieved a better result——Most SC. This is also a great example that highlights our model’s ability to collaborate effectively with other players.
@@ B.2 Case of Avoiding Deception
As shown in Figure  7, Richelieu controls Germany. During the negotiation phase, England proposed a ceasefire to Germany and invited Germany to form an alliance to attack France jointly. England hoped to cease the war with Germany in Holland and Belgium. Subsequently, German units supported England in attacking Brest, and then England utilized its fleets to assist Germany in attacking Spain and Portugal. Richelieu suspected that England was deceiving Germany, as England was likely to attack territories in the north such as Belgium and Berlin after German units were redirected to support Brest. Therefore, we pretended to accept England’s alliance proposal during the negotiation process. However, at the same time, we sought out France and expressed our willingness to cease hostilities, allowing France to focus entirely on defending against England’s attacks. In the action phase, England’s actions confirmed Richelieu’s suspicions. England attacked Belgium from Holland, but because Richelieu didn’t move units in Belgium, England’s attack failed.
@@ Appendix D Ethical Consideration
The method proposed in this work has the potential for positive uses like enabling AI agents to emerge in cooperation via negotiation or avoiding being fooled by fake promises (or helping humans do so). However, negative cases can also arise if the technique is used for possible fraud activities. Fortunately, there is research Bakhtin et al. [2019]Zellers et al. [2019] dealing with such scenarios. And we also urge for more research efforts in this field to foster safe applications of similar technologies.
@@ 8. Answer: [Yes]
  9. Justification: The paper discussed the limitations of the work performed by the authors in the section "Limitation and Future Work".
@@ 13. Answer: [Yes]
  14. Justification: The paper provides the full set of assumptions and a complete (and correct) proof for each theoretical result in the "Method" and "Experiment" sections.
@@ 18. Answer: [Yes]
  19. Justification: The paper fully discloses all the information needed to reproduce the main experimental results in the main text section "Experiment" and appendix section "Implementation Details".
@@ 23. Answer: [Yes]
  24. Justification: The paper provides open access to the data and code in the section "Experiment".
@@ 28. Answer: [Yes]
  29. Justification: The paper specify all the training and test details to train the model in the "Experiment" section and "Implementation Details" section of appendix.
@@ 33. Answer: [Yes]
  34. Justification: The paper reports error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiment in the "Experiment" section.
@@ 48. Answer: [Yes]
  49. Justification: The paper discuss both potential positive societal impacts and negative societal impacts of the work performed in Appendix section "Ethical Consideration".
@@ 53. Answer: [Yes]
  54. Justification: The paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse in Appendix section "Ethical Consideration".
@@ 63. Answer: [Yes]
  64. Justification: new assets introduced in the paper are well documented and the documentation is provided in section "Experiment".
@@ 68. Answer: [Yes]
  69. Justification: our work does not involve crowdsourcing or research with human subjects.
@@ 73. Answer: [Yes]
  74. Justification: our work does not involve research with human subjects.